OpenStudy (anonymous):

Which is the equation of the given line? The points to create the line are: (-5,-2) and (2,-2) A. y= -2 B. y= -2x C. x= -2 D. x= 2 Please explain your answer! Thanks for helping!

OpenStudy (anonymous):

m = change in y / change in x m = (-2 - (-2)) / (-5 - 2) m = (0) / (-7) m = 0 (y - b) = m (x - a) (y - (-2)) = 0 (x - 2) y + 2 = 0 y = -2 *where m := the gradient of a line and (a,b)

OpenStudy (anonymous):

The final answer is A) y = -2

OpenStudy (anonymous):

@livias.random Oh, thanks for your help!

OpenStudy (anonymous):

You're welcome!